OpenStudy (anonymous):

Dale is making a small table in the shape of a right triangle. If two of the sides will measure about 12 inches each, how many inches should the longest side (c) measure? Use the appropriate formula and round your answer to the nearest tenth. Answer 4.9 in 6.9 in 17.0 in 24.0 in

OpenStudy (anonymous):

@iNeedTheHelp

OpenStudy (anonymous):

C^2 = A^2 + B^2 Root ( C^2 ) 12^2 + 12^2 = 288 (C^2) 16.970...=17.0 option C

OpenStudy (anonymous):

24.0 in

OpenStudy (anonymous):

@iNeedTheHelp I like yours.